OpenStudy (anonymous):

a roof rises 2.3 feet vertically, and 9.5 feet horizontally. What is the grade of the roof?

OpenStudy (anonymous):

Grade = rise/run (or vertical/horizontal) Grade = 2.3/9.5 = 0.24